You have been assigned the task of using the corporate valuation model to estimate Hindelang Co.’s intrinsic value. Hindelang’s WACC is 12.00%, its expected end-of-year free cash flow (FCF1) is $150 million, the FCFs are expected to grow at a constant rate of 4.00% a year in the future, the company has $175 million of long-term debt plus preferred stock, and it has 50 million shares of common stock outstanding. What is the per-share estimate of the stock’s intrinsic value?
Answer #1

Enterprise Value=Expected Free Cash Flow/(WACC-growth rate)=150/(12%-4%)=1875 million

Equity Value=Enterprise Value-Debt-Preferred Stock=1875-175=1700 million

Price per share=Equity Value/Number of Shares=1700/50=34